# Trajectory Consistency Distillation

### v1

@@ 5.2 Main Results
For the quantitative evaluation, we conducted zero-shot image generation using 5K captions from COCO2017 validation set. We employed the Frechet Inception Distance (FID) and the Image Complexity Score (Feng et al., 2022) to assess the sample quality and visual complexity of generated images. Additionally, we used Image Reward (Xu et al., 2023a) and PickScore (Kirstain et al., 2023) to quantify text-image alignment and human preference. As depicted in Table 1, TCD shows an improvement in the performance of LCM across various timesteps and metrics. In comparison to numerical ODE-Solver, TCD exhibits the ability to generate high-quality images in only 2 or 4 steps, surpassing them significantly. When increasing steps to 8 or 20, LCM’s performance tends to decline, while TCD even outperforms its teacher in almost all metrics.
@@ 5.3 Ablation Studies
We illustrate the impact of our proposed different parameterisation types in Table 2 with 4 sampling steps. We observed that the instability of the high-order solver discussed in (Lu et al., 2022b) also exists in TCF(2), resulting in slightly inferior performance compared to TCF(1). Due to the introduction of additional parameters, TCF(S+) faces challenges in effectively distilling the teacher’s information, leading to poorer performance.
@@
To assess the versatility of TCD, we extensively tested it on various models, including the popular community model Animagine XL V3  [^1] Animagine: https://civitai.com/models/260267/animagine-xl-v3, styled LoRA Papercut XL  [^2] Papercut: https://civitai.com/models/122567/papercut-sdxl, Depth ControlNet  [^3] Depth ControlNet: https://huggingface.co/diffusers/controlnet-depth-sdxl-1.0, Canny ControlNet  [^4] Canny ControlNet: https://huggingface.co/diffusers/controlnet-canny-sdxl-1.0, and IP-Adapter  [^5] IP-Adapter: https://github.com/tencent-ailab/IP-Adapter. The results shown in Figure 7 imply that TCD can be directly applied to various models to accelerate image generation with high quality in only 2-8 steps. Additional samples can be found in Section E.4.
@@ 6 Conclusion
In this work, we introduce TCD, a novel distillation method that reduces inherent errors present in consistency models, including TCF for training and SSS for sampling. The TCF is proposed to diminish distillation errors and enable the model to track the trajectory along the PF ODE. Furthermore, SSS is proposed to reduce accumulated errors by the bijective traversal. Remarkably, TCD outperforms LCM across all sampling steps and exhibits superior performance compared to numerical methods of teacher model. We believe that TCD can provide novel perspectives for fast and high-quality image generation, while certain characters of TCD also contribute valuable insights to downstream applications, e.g., enhanced details for super-resolution and a better intermediate manifold for editing.
@@
Impact Statements
While our advancements in sample quality and speed can help reduce inference costs, they also have the potential to amplify negative societal effects, such as disseminating disinformation. In the future, it is encouraged to implement adequate guardrails and detection techniques to minimize the risk of misuse.
Appendix A Related Works
Diffusion Models.
Diffusion Models (DMs), also known as Score-based generative models, have shown superior performance in various generative fields. Chen et al. (2022) provides theoretical convergence guarantees, implying that DMs can efficiently sample from essentially any realistic data distribution under minimal data assumptions. ADM (Dhariwal & Nichol, 2021) firstly shows the potential of DMs to outperform GANs. EDM (Karras et al., 2022) further elucidates the design space of GMs, clearly separating the concrete design choices, deriving best practices for the sampling process, and improving the training dynamics, thereby drastically improving the results.
Text-conditional Diffusion Models.
Fast Sampling of DMs.
DMs exhibit great generative capabilities but are bottlenecked by their slow sampling speed. Various numerical methods driven by the SDE mathematical underpinnings underlying DMs have been proposed for accelerating DM sampling. DDIM (Song et al., 2020a) has originally shown promise for few-step sampling. Other works involve predictor-corrector samplers (Song et al., 2020b; Karras et al., 2022), exponential integrators (Lu et al., 2022a; Zhang & Chen, 2022; Lu et al., 2022b), and automated methods to adjust solvers (Kong & Ping, 2021; Jolicoeur-Martineau et al., 2021). Despite these methods having achieved great improvement in fast sampling, they are limited by the inherent discretisation error present in all solvers (De Bortoli et al., 2021). Ultimately, the sample quality obtained with few NFEs is limited. Another series of research works exemplified by distillation techniques (Luhman & Luhman, 2021; Salimans & Ho, 2022; Meng et al., 2023). These methods distill knowledge from pretrained models into few-step samplers, representing an efficient solution for few NFEs. However, they may experience a lengthy and costly process that requires huge amounts of data and suffer from slow convergence.
Consistency Models.
To overcome the limitations of current fast samplers, Song et al. (2023) proposes Consistency Models (CMs) learning a direct mapping from noise to data built on the top of the trajectory of PF ODE, achieving one-step generation while allowing multi-step sampling to trade compute for quality. Lyu et al. (2023) provide the first convergence guarantees for CMs under moderate assumptions. Kim et al. (2023) proposes a universal framework for CMs and DMs. The core design is similar to ours, with the main differences being that we focus on reducing error in CMs, subtly leverage the semi-linear structure of the PF ODE for parameterization, and avoid the need for adversarial training. Latent Consistency Models (LCMs) (Luo et al., 2023a) integrate consistency distillation with latent diffusion models (Rombach et al., 2022) to achieve impressive performance in accelerating text-to-image synthesis. LCM-LoRA (Luo et al., 2023b) further improves the training efficiency and versatility of LCMs by introducing LoRA (Hu et al., 2021) into the distillation process. However, due to inherent accumulated errors in multistep consistency samplings, LCM experiences a notable decline in quality at high NFEs. In contrast, our TCD leverages trajectory consistency and bijective traversal to eliminate this defect.
Parameter-Efficient Fine-Tuning.
@@ B.1 Algorithm Details of Trajectory Consistency Distillation
Due to space limitations, we omitted some implementation details in the main body, but we provided a detailed algorithm for trajectory consistency distillation during training as Algorithm 1 and a guided version for conditional models as Algorithm 2.
@@
     $\hat{{\bm{x}}}_{t_{n}}^{\phi,k}\leftarrow(1+\omega)\Phi^{(k)}({\bm{x}}_{t_{n+k}},t_{n+k},t_{n},{\bm{c}};\bm{\phi})-$ $\omega\Phi^{(k)}({\bm{x}}_{t_{n+k}},t_{n+k},t_{n},{\bm{c}};\bm{\phi})$
@@ B.2 The Difference between Multistep Consistency Sampling and Strategic Stochastic Sampling
     $\tilde{{\bm{x}}}_{\tau_{(n+1)}}\leftarrow\alpha_{\tau_{(n+1)}}{\bm{x}}_{\tau_{n}\rightarrow\tau_{n}^{{}^{\prime}}}+\sigma_{\tau_{(n+1)}}{\bm{z}}$
@@
     ${\bm{x}}_{\tau_{n}\rightarrow\tau_{n}^{{}^{\prime}}}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}(\tilde{{\bm{x}}}_{\tau_{n}},\tau_{n},\tau_{n}^{{}^{\prime}})$ Diffuse
@@
  ${\bm{x}}_{\tau_{N}\rightarrow 0}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}(\tilde{{\bm{x}}}_{\tau_{N}},\tau_{N},0)$

### v2

@@ 5.2 Main Results
For the quantitative evaluation, we conducted zero-shot image generation using 5K captions from COCO2017 validation set. We employed the Frechet Inception Distance (FID) and the Image Complexity Score (Feng et al., 2022) to assess the sample quality and visual complexity of generated images. Additionally, we used Image Reward (Xu et al., 2023) and PickScore (Kirstain et al., 2023) to quantify text-image alignment and human preference. As depicted in Table 1, TCD shows an improvement in the performance of LCM across various timesteps and metrics. In comparison to numerical ODE-Solver, TCD exhibits the ability to generate high-quality images in only 2 or 4 steps, surpassing them significantly. When increasing steps to 8 or 20, LCM’s performance tends to decline, while TCD even outperforms its teacher in almost all metrics.
@@ 5.3 Ablation Studies
We illustrate the impact of our proposed different parameterisation types in Table 2 with 4 sampling steps. We observed that the instability of the high-order solver discussed in (Lu et al., 2022b) also exists in TCF(2), resulting in slightly inferior performance compared to TCF(1). Due to the introduction of additional parameters, TCF(S+) faces challenges in effectively distilling the teacher’s information, leading to poorer performance. Note that while TCF(1) only requires 3000 training iterations to achieve satisfactory results in Table 2, TCF(S+) takes 43,000 iterations but still lags behind in performance.
@@
To assess the versatility of TCD, we extensively tested it on various models, including the popular community model Animagine XL V3  [^2] Animagine: https://civitai.com/models/260267/animagine-xl-v3, styled LoRA Papercut XL  [^3] Papercut: https://civitai.com/models/122567/papercut-sdxl, Depth ControlNet  [^4] Depth ControlNet: https://huggingface.co/diffusers/controlnet-depth-sdxl-1.0, Canny ControlNet  [^5] Canny ControlNet: https://huggingface.co/diffusers/controlnet-canny-sdxl-1.0, and IP-Adapter  [^6] IP-Adapter: https://github.com/tencent-ailab/IP-Adapter. The results shown in Figure 7 imply that TCD can be directly applied to various models to accelerate image generation with high quality in only 2-8 steps. Additional samples can be found in Section D.4.
@@ 6 Conclusion
In this work, we introduce TCD, a improved latent consistency distillation method that reduces inherent errors present in latent consistency models, including TCF for training and SSS for sampling. The TCF is proposed to diminish parameterisation and distillation errors and enable the model to track the trajectory along the PF ODE with a semi-linear structure. Furthermore, SSS is proposed to explicitly control the stochastic intensity and reduce accumulated errors by the bijective traversal. Remarkably, TCD outperforms LCM across all sampling steps and exhibits superior performance compared to numerical methods of teacher model. We believe that TCD can provide novel perspectives for fast and high-quality image generation, while certain characters of TCD also contribute valuable insights to downstream applications, e.g., enhanced details for super-resolution and a better intermediate manifold for editing.
@@
Appendix A Discussion between CTM (Kim et al., 2023) and TCD
A recent work, CTM (Kim et al., 2023), proposed a unified framework for CM and DM solvers, with the idea of “anytime to anytime” prediction, theoretically substantiated and facilitated through a $G_{\theta}$ function with $\gamma$-sampling for assessing both the integrand and the integral of PF-ODE. Conversely, TCD focused on improving the performance of LCM by leveraging a semi-linear structure with a narrower time spacing in the consistency function. While these two studies share intersections in their approaches, such as the use of varying time intervals from $t$ to $s$, they also exhibit distinct differences. In this section, we delve into a discussion on CTM and TCD, exploring their various aspects.
To initiate our comparison, we first recall two definitions: explicit numerical method and implicit numerical method for ODEs (Atkinson et al., 2009; Süli, 2010):
A general explicit method can be:
$y_{n+1}=y_{n}+hf(y_{n},t_{n}),\,\,n=0,\dots,N-1,$ (36)
where $f(y_{n},t_{n})$ is a continuous function and $h$ is the interval between $t_{n}$ and $t_{n+1}$. In an explicit method, $y_{n}$ is already calculated, and $y_{n+1}$ can be given directly.
Similarly, the implicit method, which is also said to be absolutely stable, can be expressed in the form:
$y_{n+1}=y_{n}+hf(y_{n+1},t_{n+1}),\,\,n=0,\dots,N-1,$ (37)
which requires the solution of an implicit equation $f(y_{n+1},t_{n+1})$ in order to determine $y_{n+1}$.
A representative example of an explicit method is the Euler Method, while the Backward Euler Method is an implicit method. Informally, score-based sampling with time-discretized numerical integral solvers can be regarded as an explicit method, whereas the distillation model that directly estimates the integral of ODEs is an implicit method (Kim et al., 2023).
In order to unify CMs and DMs by assessing both the infinitesimal jump and long jump, CTM (Kim et al., 2023) proposed a unified decoder $G$ with an “anytime-to-anytime” idea, where $G$ is the solution of the PF-ODE with an integrator term. Furthermore, CTM used a trainable $g_{\theta}$ to parameterize the integrator term and made it access the infinitesimal jump inspired by the Euler Solver. In order to train such a $G_{\theta}$ function, instead of directly forcing $G_{\theta}({\bm{x}}_{t},t,s)\approx\texttt{Solver}({\bm{x}}_{t},t,s;\phi)$, CTM proposed a soft consistency matching $G_{\theta}({\bm{x}}_{t},t,s)\approx G_{\text{sg}(\theta)}(\texttt{Solver}({\bm{x}}_{t},t,u;\phi),u,s)$ that supports both $u=s$ and $u=t-\Delta t$. It can be observed that this loss is similar to the distillation loss for consistency model (Song et al., 2020a). Additionally, CTM has also introduced DSM loss and GAN loss to optimize the final performance of the model. It’s evident that the adversarial loss plays an important role in reaching a better FID. To further leverage the implicit method with a well-trained $g_{\theta}$ function, CTM proposed $\gamma$-sampling, which integrates the properties of both the EDM sampler and the deterministic ODE sampler, and falls within the protocol of multistep consistency sampling, i.e., iterative denoising and diffusing. If we were to apply the informal analogy of explicit and implicit methods, CTM proposed a distillation-based implicit method with $G_{\theta}$ to address the problems in time-discretized explicit solvers.
In contrast, TCD aims to tackle the challenge posed by the imprecise parameterisation of the consistency function (Song et al., 2023) in LCM (Luo et al., 2023a) and improve the generation quality on text-to-image task. Our method strictly adheres to the definitions of the consistency model, enabling the direct adoption of characteristics and algorithms inherent to consistency models. We first observed that several errors in the distillation process are related to the time interval $t\rightarrow s$ of the consistency function. Thus, we leverage the semi-linear structure with exponential integrators of the PF-ODE for parameterization, which also supports a shorter interval (i.e., moderately moving the upper limit $s$ away from 00). Such concept of transitioning from $t$ to $s$ overlaps with CTM’s “anytime to anytime” idea but ours is from different view. Moreover, this semi-linear structure supports ${\bm{x}}_{0}$, ${\bm{v}}$, and $\epsilon$ parameterisation, allowing for efficient distillation from large-scale diffusion models. During training, we simply follow all the definitions and targets of Consistency Distillation and equip $\mathbb{E}[\omega(t_{n},t_{m})\|{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+k}},t_{n+k},t_{m})-{\bm{f}}_{\bm{\theta}^{-}}(\hat{{\bm{x}}}_{t_{n}}^{\phi,k},t_{n},t_{m})\|^{2}_{2}]$ as the loss while $\hat{{\bm{x}}}_{t_{n}}^{\phi,k}$ is from a solver by 1 NFE with parameter $\phi$ and $k\geq 1$. As observed in (Sauer et al., 2023), GAN loss tends to result in poorer mode coverage, hence we only employ the distillation loss used in (Song et al., 2023; Luo et al., 2023a) to ensure the diversity of generated images. The sampling process of TCD remains explicit as the TCF can be easily parameterized by $\epsilon$, allowing it to directly utilize existing explicit samplers. This means that once trained, TCF functions as a plug-and-play model, eliminating the need for designing additional samplers. Recognizing Multistep Consistency Sampling with parameterisation consistency function in LCM as a fixed stochasticity DDIM solver (Equation 31), we introduced SSS to achieve a flexible stochastic intensity control. Since CTM (Kim et al., 2023) proposed a unified framework for CMs and DMs and our SSS is directly based on multistep sampling in CMs, it demonstrates theoretical similarity to the $\gamma$-sampling used in CTM (Kim et al., 2023). And the theoretical insights provided by CTM (Kim et al., 2023) in sampling errors can also support our conjectures, e.g., injecting more noise will lead to degraded quality.
In summary, CTM has four key contributions: 1. A single model inspired by Euler Solver that attains all the advantages of score-based models and distillation models ; 2. $\gamma$-sampling for flexible sampling; 3. Soft matching for efficient and stabilized training; 4. The use of auxiliary losses for better student training. Meanwhile, TCD is an enhancement of LCM (Luo et al., 2023a), adhering to the definition of consistency models with the following features: 1. Better parameterization of consistency functions with a semi-linear structure (Lu et al., 2022a, b; Zhang & Chen, 2022); 2. Maintains the explicit sampling method and is versatile for existing solvers; 3. Employs only consistency loss with skipping-steps to achieve self-consistency and avoid mode collapse. Moreover, TCD aims to improve the quality of general generation tasks in fewer steps, instead of achieving SOTA for single-step generation for specific tasks.
@@ B.1 Algorithm Details of Trajectory Consistency Distillation
We provided a detailed algorithm for trajectory consistency distillation during training as Algorithm 1 and a guided version for conditional models as Algorithm 2.
@@
     $\hat{{\bm{x}}}_{t_{n}}^{\phi,k}\leftarrow(1+\omega)\Phi^{(k)}({\bm{x}}_{t_{n+k}},t_{n+k},t_{n},{\bm{c}};\bm{\phi})-$ $\omega\Phi^{(k)}({\bm{x}}_{t_{n+k}},t_{n+k},t_{n},\varnothing;\bm{\phi})$
@@ B.2 The Difference between Multistep Consistency Sampling and Strategic Stochastic Sampling
     $\hat{{\bm{x}}}_{\tau_{(n+1)}}\leftarrow\alpha_{\tau_{(n+1)}}{\bm{x}}_{\tau_{n}\rightarrow\tau_{n}^{{}^{\prime}}}+\sigma_{\tau_{(n+1)}}{\bm{z}}$
@@
     ${\bm{x}}_{\tau_{n}\rightarrow\tau_{n}^{{}^{\prime}}}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}^{\rightarrow\tau_{n}^{{}^{\prime}}}(\hat{{\bm{x}}}_{\tau_{n}},\tau_{n})$ Diffuse
@@
  ${\bm{x}}_{\tau_{N}\rightarrow 0}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}^{\rightarrow 0}(\tilde{{\bm{x}}}_{\tau_{N}},\tau_{N})$